OpenStudy (anonymous):

Mr Castle Shares out 64 marbles Emily and Asif in the ratio 7:1How many Marbles does each child get? Emily gets...... Marbles and Asif Gets.....Marbles

OpenStudy (anonymous):

Coud U Help Plz

OpenStudy (anonymous):

you gotta divide 64 in 7 :1

OpenStudy (anonymous):

kl

OpenStudy (anonymous):

let say x then total number is 7x+1=64 7x=64-1=63 x=9

OpenStudy (anonymous):

so wats the answer

OpenStudy (anonymous):

how much does emily get and how much does asif get

OpenStudy (anonymous):

Plz can u tell me the answer

OpenStudy (anonymous):

callistor could you help me how much does emily get and how much does asif get

OpenStudy (callisto):

''Mr Castle Shares out 64 marbles Emily and Asif in the ratio 7:1'' in the ratio of 7:1 => when there are 8 marbles, Emily will get 7, Asif will get 1 Now, there are 64 marbles, so, what do you think?

OpenStudy (anonymous):

I Dont Know

OpenStudy (callisto):

Do you understand ''in the ratio of 7:1 => when there are 8 marbles, Emily will get 7, Asif will get 1'' ?

OpenStudy (anonymous):

I dont Know could u just tell me how much emily gets and how much asif gets

OpenStudy (anonymous):

plzzzzzzzzzzzzzzzzzzzzzzz